# Sard properties for polynomial maps in infinite dimension

### v2

@@ 1.1. The Sard property for polynomial maps on infinite-dimensional spaces
The classical Sard theorem [Mor39, Sar42] states that the set of critical values of a smooth map $f:N\to M$, where $N$ and $M$ are finite dimensional manifolds, has measure zero in $M$. Smale [Sma65] proved a version of this result in the case both $M$ and $N$ are infinite dimensional Banach manifolds and the map $f$ is Fredholm: in this case the conclusion is that the set of regular values of $f$ is residual in $N$.
When $N$ is infinite-dimensional and $M$ is finite dimensional, $f$ cannot be Fredholm and Smale’s result cannot be applied. In fact there are smooth surjective maps without regular points from any infinite-dimensional Banach space to $\mathbb{R}^{2}$, see [Bat93]. Even under the hypothesis that $f$ is a polynomial map, the set of its critical values can be of positive measure: Kupka [Kup65] constructed an example of a smooth map $f:\ell^{2}\to\mathbb{R}$, whose restriction to each finite dimensional space is a polynomial of degree $3$, and with the property that the set of its critical values is the segment $[0,1]$. On the other hand, using a quantitative version of the classical Sard theorem [Yom83], Yomdin has shown that if a map $f:\ell^{2}\to\mathbb{R}$ can be approximated well-enough with finite dimensional polynomials, still the set of its critical values has measure zero, see [Yom88]. (Kupka’s map does not have this property.) In Yomdin’s proof, the fact that the codomain is one-dimensional is essential and new technical difficulties arise for maps with values in $\mathbb{R}^{m}$, $m>1$.
@@
The map constructed by Kupka belongs to this family, as well as the Endpoint maps for horizontal path spaces on Carnot groups (see Section 1.3).
We need to introduce a notion of quantitative compactness for subsets of Hilbert spaces. More precisely, given $K\subset H$ and $n\in\mathbb{N}$, we denote by $\Omega_{n}(K,H)$ its Kolmogorov $n$–width:
@@
[…]
Our results are stated in terms of entropy dimension of the set of $\nu$-critical values of a map $f:H\to\mathbb{R}^{m}$, namely $f(\operatorname{Crit}_{\nu}(f))$, where $\operatorname{Crit}_{\nu}(f)$ is the set of points where $Df$ has rank at most $\nu\leq m-1$. The classical set of critical values corresponds to setting $\nu=m-1$, and it is denoted by $f(\operatorname{Crit}(f))$, where $\operatorname{Crit}(f)$ is the set of critical points. We refer to Definition 24 of entropy dimension (denoted by $\dim_{e}$, and also known as Minkowski dimension), noting that it is larger than the Hausdorff one so that in the forthcoming estimates one can replace the former with the latter for simplicity.
Our first theorem is a sufficient condition for the validity of the Sard property (see Theorem 31). We denote by $\mu$ the Lebesgue measure on $\mathbb{R}^{m}$.
Theorem A (Sard under $n$–width assumptions).
@@
The constant $\beta_{0}(d,m)$ of Theorem A comes from semialgebraic geometry, and its origin is discussed in Section 1.2. For this reason, we sometimes call it the semialgebraic constant.
By building upon Kupka’s counterexamples [Kup65], we prove that the quantitative assumption $q>\beta_{0}$ in Theorem A cannot be dispensed of. The following result corresponds to Theorem 35. We denote by $B_{X}(r)$ the closed ball of radius $r$ in a Banach space $X$.
@@
and $f:H\to\mathbb{R}^{m}$ does not verify the Sard property, namely $\mu(f(\operatorname{Crit}(f)\cap K))>0$. Therefore, the semialgebraic constant $\beta_{0}(d,m)$ of Theorem A satisfies
@@
The case $d=2$, which is left out from Theorem B, is special. This is related to the fact that the complexity of semialgebraic sets defined by quadratic equations has a different behaviour compared to the case $d\geq 3$. In this case we currently expect that any $f\in\mathscr{P}^{m}_{2}(H)$ satisfies the Sard property globally on $H$. This will be the subject of a future work.
Replacing the entropy dimension with the Hausdorff dimension (which is smaller than the entropy dimension, but countably stable) we strengthen Theorem A as follows (see Theorem 37).
@@
Let $d,m\in\mathbb{N}$, let $\beta_{0}=\beta_{0}(d,m)>1$ be the same constant of Theorem A. Consider the (possibly non-closed) linear subspace
@@
In all previous results, $f\big{(}\mathrm{Crit}(f|_{V})\big{)}$ can be replaced with the smaller set $f(\mathrm{Crit}(f)\cap V)$.
@@
Let $q\in(1,\infty)$ with $q>\beta_{0}(d,m)$, where $\beta_{0}(d,m)>1$ is the number in Theorems A and C, for $d,m\in\mathbb{N}$. Let $H=\ell^{2}$ with the usual norm. Consider the subset $K\subset B_{H}(1)$ given by
@@
One can easily verify that $K$ is compact. Furthermore, the linear space $V=\mathrm{span}(K)$ is dense in $H$. We can estimate the $n$–width of $K$ as follows. For $n\in\mathbb{N}$ consider the $n$-dimensional subspaces $E_{n}=\{x\in\ell^{2}\mid x_{j}=0,\,\forall\,j\geq n+1\}\subset H$. From (2), we obtain
@@
Hence $\limsup_{n}\Omega_{n}(K,H)^{1/n}\leq q^{-1}$, with $q>\beta_{0}(d,m)$. Thus, by Theorem C it holds (in particular)
@@
Note that even if $V\subset H$ is dense, the unrestricted map $f:H\to\mathbb{R}^{m}$ may not have the Sard property (see e.g. the Kupka counterexamples in Section 4.3).
Theorems A, F and C are deduced from a more general result for maps that are “well-approximated” by polynomials. This can be regarded as our main result concerning the Sard property, and does not make use of the concept of $n$–width. We report here the statement (see Theorem 25).
@@
Let $d,m\in\mathbb{N}$. There exists a constant $\beta_{0}=\beta_{0}(d,m)>1$ such that the following holds. Let $H$ be a Hilbert space, and let $f:H\to\mathbb{R}^{m}$ be a $\mathcal{C}^{1}$ map such that its differential $Df:H\to\mathcal{L}(H,\mathbb{R}^{m})$ is weakly continuous. Let $K\subset H$ be a bounded set with this approximation property: there exist a sequence $E_{n}\subset H$ of linear subspaces, $\dim(E_{n})=n$, and polynomial maps $f_{n}:E_{n}\to\mathbb{R}^{m}$ with uniformly bounded degree:
@@
(18) $\sup_{x\in K}\bigg{(}\|f(x)-f_{n}\circ\pi_{E_{n}}(x)\|+\|\left(D_{x}f\right)|_{E_{n}}-D_{\pi_{E_{n}}(x)}f_{n}\|_{\operatorname{op}}\bigg{)}\leq cq^{-n}.$
@@
In particular, if $q>\beta_{0}$, then $f$ satisfies the Sard property on $K$:
(20) $\mu\bigg{(}f(\operatorname{Crit}(f)\cap K)\bigg{)}=0.$
By means of the general Theorem D we single out a class of maps satisfying the Sard property on the whole domain of definition. This is the content of Theorem 41, that we report here. The case $m=1$ corresponds to [YC04, Thm. 10.12].
@@
is well–defined, $f\in\mathscr{P}_{d}^{m}(H)$ (see Definition 1), and for all $\nu\leq m-1$ and $r>0$ it holds
@@
where $\beta_{0}=\beta_{0}(d,m)>1$ is the same constant given by Theorem D. In particular, if $q>\beta_{0}$, then $f$ satisfies the Sard property globally on $H$:
@@
As a consequence of Theorems A and B, we see that there is a threshold on the $n$–width for the validity of the Sard property. Motivated by this fact, for a compact set $K\subset H$ we define the quantity:
@@
Roughly speaking, smaller $\omega(K,H)$ means that $K$ is “more compact” in $H$, and better approximated by finite-dimensional subspaces. We obtain the following statement (see Theorem 42).
@@
By construction, and Theorem A, $\omega_{0}\geq\beta_{0}^{-1}$. Furthermore, if $d\geq 3$, Theorem B yields $\omega_{0}<1$, so that Item ii is non-vacuous in these cases.
@@ 1.2. Quantitative semialgebraic geometry
The proof of Theorem D needs some fine properties of semialgebraic sets. The strategy of the proof is conceptually similar to Yomdin’s [Yom88], and uses the theory of variations, introduced by Vitushkin [Vit55, Vit61] and developed in [YC04].
Let us denote with $V_{i}(S)$ the $i$-th variation of a semialgebraic set, which is a sort of $i$-dimensional volume, (see Definition 14). Furthermore, recall that for a $C^{1}$ map $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$, and $\Lambda=(\Lambda_{1},\ldots,\Lambda_{m})\in\mathbb{R}^{m}_{+}$ the set of almost-critical values of $F$ is defined by
@@
where $\sigma_{1}(D_{x}F)\geq\cdots\geq\sigma_{m}(D_{x}F)$ are the singular values of $D_{x}F$, see Definition 16 (here we assume $n\geq m$). In [YC04, Cor. 7.4] a quantitative estimate on the variations of the almost-critical values of polynomial maps has been obtained. In that estimate the dimensional parameter $n$ does not appear explicitly. We obtain Theorem G below, which makes this dependence explicit (see Theorem 23).
@@
A key technical step in the proof of Theorem G is a quantitative substitute for the definable choice theorem in semialgebraic geometry, which we prove in [LRT24] (see Theorem 13).
@@ 1.3. Applications to Endpoint maps of Carnot groups
We discuss now the implications of Theorem F and Theorem C in the context of sub-Riemannian geometry.
@@
where $\mathfrak{g}_{i}\neq\{0\}$, $\mathfrak{g}_{i+1}=[\mathfrak{g}_{1},\mathfrak{g}_{i}]$ for all $i=1,\dots,s-1$ and $[\mathfrak{g}_{1},\mathfrak{g}_{s}]=\{0\}$. The group exponential map yields an identification $\mathbb{G}\simeq\mathbb{R}^{m}$, and the first stratum $\mathfrak{g}_{1}$ of the Lie algebra defines a smooth, totally non-holonomic distribution $\Delta\subseteq T\mathbb{R}^{m}$ of rank $k:=\dim\mathfrak{g}_{1}$.
@@
The conjecture was introduced by Zhitomirskii and Montgomery, see [Mon02, Sec. 10.2] for general sub-Riemannian structures, where the Endpoint map can be defined in a similar way, on a suitable domain of horizontal paths, which has the structure of a Hilbert manifold. For the specific case of Carnot groups, it is known to be true for step $\leq 2$, see [AGL15, LDMO+16]. Furthermore, the conjecture has been verified in [BNV22] for filiform Carnot groups, in [BV20] for Carnot groups of rank 2 and step 4, and rank 3 and step 3, and for a handful of other specific examples described in [LDMO+16].
Assume that $\Delta$ is also endowed with a left-invariant norm. This choice induces a length structure on the space of horizontal curves, and correspondingly, a distance, which gives $\mathbb{G}$ the structure of sub-Finsler Carnot group (or sub-Riemannian Carnot group if the norm is induced by a scalar product). In this case, within the set critical points of $\mathcal{E}$ (also called singular horizontal paths), the energy-minimizing ones play a significant role. At the corresponding critical values, the distance loses regularity: it is never smooth and it can even lose local semiconcavity [BR19, Sec. 4.2]. For this reason, the understanding of the “abundance” of such critical values is crucial. Denoting with $H_{\min}\subseteq L^{2}(I,\mathbb{R}^{k})$ the set of minimizing horizontal paths starting from the identity, the minimizing Sard conjecture can be formulated as follows:
@@
The most general result to date is that the above set is a closed nowhere dense set [Agr09, RT05], which of course does not imply the minimizing Sard conjecture. The above conjecture is one of the main open problems in sub-Riemannian geometry [Agr14, Prob. 3], [RT05, Conj. 1]. In general, the problem is settled in the following cases:
    Thanks to the work of Agrachev-Sarychev [AS96] and Agrachev-Lee [AL09], in absence of so-called Goh singular minimizing paths, all sub-Riemannian minimizing paths (resp. sub-Finsler, under suitable smoothness assumptions on the norm) are solution of a finite-dimensional Hamiltonian flow so that the corresponding set of critical values has zero measure by the finite-dimensional Sard theorem. Recently, Rifford proved that, more in general, the minimizing Sard conjecture holds for structures where all non-trivial Goh paths have Goh-rank $\leq 1$ almost everywhere, see [Rif23]. For Carnot groups, these results can be applied for example when the distribution is pre-medium fat, see [Rif23, (1.2)]. In passing, we remark that for generic (so, typically not Carnot) sub-Riemannian structures with distribution of rank $\geq 3$, there are no non-trivial singular Goh minimizing paths, see [CJT06, Cor. 2.5] and [AG01, Thm. 8]. As a consequence the minimizing Sard conjecture holds true. Thanks to [Rif23, Cor. 1.4], the latter result extends the case of rank $\geq 2$.
    The minimizing Sard conjecture is true also for Carnot groups of step $\leq 3$, see [LDMO+16]. In this case, the problem is reduced (in a non-trivial way) to a finite-dimensional one. First, by noting that singular minimizing curves are non-singular in some proper subgroup, and then exploiting the fact that Carnot subgroups are parametrized by a finite-dimension manifold.
Finally, we refer to [ZZ95, BdSR18, BdSFPR22, BdSPR23, BdSPR22, OV19, RT05] for further details and discussions of various forms of the Sard conjecture on general sub-Riemannian structures.
The following result (Theorem 46) connects this framework to the previous sections.
Theorem H (Polynomial properties of the Endpoint map).
@@
As a consequence of Theorem H, we can apply Theorem C to the study of Endpoint maps of Carnot groups. We record here a first immediate corollary.
Theorem I (Sard criterion for Endpoint maps).
@@
where $\beta_{0}=\beta_{0}(s,m)>1$ is the same constant of Theorem A. In particular, it holds:
@@
We can apply Theorem I to find large sets on which the Sard property holds for the Endpoint map of Carnot groups. This is a more functional-analytic approach to the Sard conjecture that, unlike previous ones, does not resort to reduction to finite-dimensional cases.
An interesting case in which the assumptions of Theorem I can be effectively checked is given by sets of controls that are “sufficiently regular”. We introduce the notation to state our result, see Section 5.2 for additional details. Given $r>0$, and a closed interval $I\subset\mathbb{R}$, consider the set:
@@
endowed with the sup norm over the $r$-neighbourhood of $I$ in $\mathbb{C}$. More in general, let $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$ be the set of piecewise analytic controls:
@@
where $I_{j}=\inf I+\left[\frac{(j-1)|I|}{\ell},\frac{j|I|}{\ell}\right]$, for all $j=1,\dots,\ell$, which we equip with a suitable norm (see Section 5.2). In Theorem 47 we prove that the unit ball $K$ of $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$, with $I=[0,1]$, is compact in $H$ and its $n$–width satisfies:
@@
By applying Theorem I we can therefore deduce the following result (see Theorem 48).

### v3

@@ 1.1. The Sard property for polynomial maps on infinite-dimensional spaces
The classical Sard theorem [Morse, Sard] states that the set of critical values of a smooth map $f:N\to M$, where $N$ and $M$ are finite dimensional manifolds, has measure zero in $M$. Smale [smale] proved a version of this result in the case both $M$ and $N$ are infinite dimensional Banach manifolds and the map $f$ is Fredholm: in this case the conclusion is that the set of regular values of $f$ is residual in $N$.
When $N$ is infinite-dimensional and $M$ is finite dimensional, $f$ cannot be Fredholm and Smale’s result cannot be applied. In fact there are smooth surjective maps without regular points from any infinite-dimensional Banach space to $\mathbb{R}^{2}$, see [Bates]. Even under the hypothesis that $f$ is a polynomial map, the set of its critical values can be of positive measure: Kupka [Kupka] constructed an example of a smooth map $f:\ell^{2}\to\mathbb{R}$, whose restriction to each finite dimensional space is a polynomial of degree $3$, and with the property that the set of its critical values is the segment $[0,1]$. On the other hand, using a quantitative version of the classical Sard theorem [Yomdinnearly], Yomdin has shown that if a map $f:\ell^{2}\to\mathbb{R}$ can be approximated well-enough with finite dimensional polynomials, still the set of its critical values has measure zero, see [YomdinApproxCompl]. (Kupka’s map does not have this property.) In Yomdin’s proof, the fact that the codomain is one-dimensional is essential and new technical difficulties arise for maps with values in $\mathbb{R}^{m}$, $m>1$.
@@
The map constructed by Kupka belongs to this family, as well as the Endpoint maps for horizontal path spaces on Carnot groups (see Section˜1.3).
We need to introduce a notion of quantitative compactness for subsets of Hilbert spaces. More precisely, given $K\subset H$ and $n\in\mathbb{N}$, we denote by $\Omega_{n}(K,H)$ its Kolmogorov nn–width:
@@
where $\|\cdot\|$ denotes the norm of the Hilbert space $H$, and the infimum ranges over all vector subspaces $E\subset H$ of dimension $n$ (see Definition˜30). A set $K\subset H$ is compact if and only if it is bounded and its nn–width goes to zero as $n$ goes to infinity.
Our results are stated in terms of entropy dimension of the set of $\nu$-critical values of a map $f:H\to\mathbb{R}^{m}$, namely $f(\operatorname{Crit}_{\nu}(f))$, where $\operatorname{Crit}_{\nu}(f)$ is the set of points where $Df$ has rank at most $\nu\leq m-1$. The classical set of critical values corresponds to setting $\nu=m-1$, and it is denoted by $f(\operatorname{Crit}(f))$, where $\operatorname{Crit}(f)$ is the set of critical points. We refer to Definition˜27 of entropy dimension (denoted by $\dim_{e}$, and also known as Minkowski dimension), noting that it is larger than the Hausdorff one so that in the forthcoming estimates one can replace the former with the latter for simplicity.
Our first theorem is a sufficient condition for the validity of the Sard property (see Theorem˜35). We denote by $\mu$ the Lebesgue measure on $\mathbb{R}^{m}$.
Theorem A (Sard under nn–width assumptions).
@@
The constant $\beta_{0}(d,m)$ of Theorem˜A comes from semialgebraic geometry, and its origin is discussed in Section˜1.2. We sometimes call it the semialgebraic constant.
By building upon Kupka’s counterexamples [Kupka], we prove that the quantitative assumption $q>\beta_{0}$ in Theorem˜A cannot be dispensed of. The following result corresponds to Theorem˜39. We denote by $B_{X}(r)$ the closed ball of radius $r$ in a Banach space $X$.
@@
and $f:H\to\mathbb{R}^{m}$ does not verify the Sard property, namely $\mu(f(\operatorname{Crit}(f)\cap K))>0$. Therefore, the semialgebraic constant $\beta_{0}(d,m)$ of Theorem˜A satisfies
@@
The case $d=2$, which is left out from Theorem˜B, is special. This is related to the fact that the complexity of semialgebraic sets defined by quadratic equations has a different behaviour compared to the case $d\geq 3$. We currently expect that any $f\in\mathscr{P}^{m}_{2}(H)$ satisfies the Sard property globally on $H$. This will be the subject of a future work.
Replacing the entropy dimension with the Hausdorff one (which is smaller than the former, but countably stable) we strengthen Theorem˜A as follows (see Theorem˜42).
@@
Let $d,m\in\mathbb{N}$, let $\beta_{0}=\beta_{0}(d,m)>1$ be the same constant of Theorem˜A. Consider the (possibly non-closed) linear subspace
@@
In all previous results, $f\big(\mathrm{Crit}_{\nu}(f|_{V})\big)$ can be replaced with the smaller set $f(\mathrm{Crit}_{\nu}(f)\cap V)$.
@@
Let $q\in(1,\infty)$ with $q>\beta_{0}(d,m)$, where $\beta_{0}(d,m)>1$ is the number in Theorems˜A and C, for $d,m\in\mathbb{N}$. Let $H=\ell^{2}$ with the usual norm. Consider the subset $K\subset B_{H}(1)$ given by
@@
One can easily verify that $K$ is compact. Furthermore, the linear space $V=\mathrm{span}(K)$ is dense in $H$. We can estimate the nn–width of $K$ as follows. For $n\in\mathbb{N}$ consider the $n$-dimensional subspaces $E_{n}=\{x\in\ell^{2}\mid x_{j}=0,\,\forall\,j\geq n+1\}\subset H$. From (3), we obtain
@@
Hence $\limsup_{n}\Omega_{n}(K,H)^{1/n}\leq q^{-1}$, with $q>\beta_{0}(d,m)$. Thus, by Theorem˜C it holds (in particular)
@@
Note that even if $V\subset H$ is dense, the unrestricted map $f:H\to\mathbb{R}^{m}$ may not have the Sard property (see e.g. the Kupka counterexamples in Section˜4.3).
Theorems˜A, F and C are deduced from a more general result for maps that are “well-approximated” by polynomials. This can be regarded as our main result concerning the Sard property, and does not make use of the concept of nn–width. We report here the statement (see Theorem˜28).
@@
Let $d,m\in\mathbb{N}$. There exists a constant $\beta_{0}=\beta_{0}(d,m)>1$ such that the following holds. Let $H$ be a Hilbert space, and let $f:H\to\mathbb{R}^{m}$ be a $\mathcal{C}^{1}$ map such that its differential $Df:H\to\mathcal{L}(H,\mathbb{R}^{m})$ is weakly continuous. Let $B\subset H$ be a bounded set with this approximation property: there exist a sequence $E_{n}\subset H$ of linear subspaces, $\dim(E_{n})=n$, and polynomial maps $f_{n}:E_{n}\to\mathbb{R}^{m}$ with uniformly bounded degree:
@@
(19) $\sup_{x\in B}\bigg(\|f(x)-f_{n}\circ\pi_{E_{n}}(x)\|+\|\left(D_{x}f\right)|_{E_{n}}-D_{\pi_{E_{n}}(x)}f_{n}\|_{\operatorname{op}}\bigg)\leq cq^{-n}.$
@@
(20) $\dim_{e}\bigg(f\big(\operatorname{Crit}_{\nu}(f)\cap B\big)\bigg)\leq\nu+\frac{\ln\beta_{0}}{\ln q}.$
In particular, if $q>\beta_{0}$, then $f$ satisfies the Sard property on $B$:
(21) $\mu\bigg(f(\operatorname{Crit}(f)\cap B)\bigg)=0.$
By means of the general Theorem˜D we single out a class of maps satisfying the Sard property on the whole domain of definition. This is the content of Theorem˜46, that we report here. The case $m=1$ corresponds to [ComteYomdin, Thm. 10.12]. In the statement, we fix a Hilbert basis $\{e_{i}\}_{i\in\mathbb{N}}\subset H$. For $k\in\mathbb{N}$, we set $E_{k}:=\mathrm{span}\{e_{1},\ldots,e_{k}\}$ and we denote by $\pi_{k}:H\to E_{k}$ the corresponding orthogonal projection, that is if $x=\sum_{i}x_{i}e_{i}$, then $\pi_{k}(x)=(x_{1},\dots,x_{k})$.
@@
is well–defined, $f\in\mathscr{P}_{d}^{m}(H)$ (see Definition˜1), and for all $\nu\leq m-1$ and $r>0$ it holds
@@
where $\beta_{0}=\beta_{0}(d,m)>1$ is the same constant given by Theorem˜D. In particular, if $q>\beta_{0}$, then $f$ satisfies the Sard property globally on $H$:
@@
As a consequence of Theorems˜A and B, we see that there is a threshold on the nn–width for the validity of the Sard property. Motivated by this fact, for a compact set $K\subset H$ we define the quantity:
@@
Roughly speaking, smaller $\omega(K,H)$ means that $K$ is “more compact” in $H$, and better approximated by finite-dimensional subspaces. We obtain the following statement (see Theorem˜47).
@@
By construction, and Theorem˜A, $\omega_{0}\geq\beta_{0}^{-1}$. Furthermore, if $d\geq 3$, Theorem˜B yields $\omega_{0}<1$, so that Item˜ii is non-vacuous in these cases.
@@ 1.2. Quantitative semialgebraic geometry
The proof of Theorem˜D needs some fine properties of semialgebraic sets. The strategy of the proof is conceptually similar to Yomdin’s [YomdinApproxCompl], and uses the theory of variations, introduced by Vitushkin [Vit1, Vitbook] and developed in [ComteYomdin].
Let us denote by $V_{i}(S)$ the $i$-th variation of a semialgebraic set, which is a sort of $i$-dimensional volume, (see Definition˜15). Furthermore, recall that for a $C^{1}$ map $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$, and $\Lambda=(\Lambda_{1},\ldots,\Lambda_{m})\in\mathbb{R}^{m}_{+}$ the set of almost-critical values of $F$ is defined by
@@
where $\sigma_{1}(D_{x}F)\geq\cdots\geq\sigma_{m}(D_{x}F)$ are the singular values of $D_{x}F$, see Definition˜17 (here we assume $n\geq m$). In [ComteYomdin, Cor. 7.4] a quantitative estimate on the variations of the almost-critical values of polynomial maps has been obtained. In that estimate the dimensional parameter $n$ does not appear explicitly. We obtain Theorem˜G below, which makes this dependence explicit (see Theorem˜26).
@@
A key technical step in the proof of Theorem˜G is a quantitative substitute for the definable choice theorem in semialgebraic geometry, which we prove in [LRT-selection] (see Theorem˜14).
@@ 1.3. Applications to Endpoint maps of Carnot groups
We discuss now the implications of Theorem˜F and Theorem˜C in the context of sub-Riemannian geometry.
@@
where $\mathfrak{g}_{i}\neq\{0\}$, $\mathfrak{g}_{i+1}=[\mathfrak{g}_{1},\mathfrak{g}_{i}]$ for all $i=1,\dots,s-1$ and $\mathfrak{g}_{1},\mathfrak{g}_{s}]=\{0\}$. The group exponential map yields an identification $\mathbb{G}\simeq\mathbb{R}^{m}$, and the first stratum $\mathfrak{g}_{1}$ of the Lie algebra defines a smooth, totally non-holonomic distribution $\Delta\subseteq T\mathbb{R}^{m}$ of rank $k:=\dim\mathfrak{g}_{1}$.
@@
The conjecture was introduced by Zhitomirskii and Montgomery, see [Montgomerybook, Sec. 10.2] for general sub-Riemannian structures, where the Endpoint map can be defined in a similar way, on a suitable domain of horizontal paths, which has the structure of a Hilbert manifold. For the specific case of Carnot groups, it is known to be true for step $\leq 2$, see [AGL-pathspace, LDMOPV-Sardprop]. Furthermore, the conjecture has been verified in [BNV-Filiform] for filiform Carnot groups, in [BV-Dynamical] for Carnot groups of rank 2 and step 4, and rank 3 and step 3, and for a handful of other specific examples described in [LDMOPV-Sardprop].
Assume that $\Delta$ is also endowed with a left-invariant norm. This choice induces a length structure on the space of horizontal curves, and correspondingly, a distance, which gives $\mathbb{G}$ the structure of sub-Finsler Carnot group (or sub-Riemannian Carnot group if the norm is induced by a scalar product). In this case, within the set of critical points of $\mathcal{E}$ (also called singular horizontal paths), the energy-minimizing ones play a significant role. At the corresponding critical values, the distance loses regularity: it is never smooth and it can even lose local semiconcavity [BR-Inv, Sec. 4.2]. For this reason, the understanding of the “abundance” of such critical values is crucial. Denoting by $H_{\min}\subseteq L^{2}(I,\mathbb{R}^{k})$ the set of minimizing horizontal paths starting from the identity, the minimizing Sard conjecture can be formulated as follows:
@@
The most general result to date is that the above set is a closed nowhere dense set [agrasmoothness, RT-MorseSard], which of course does not imply the minimizing Sard conjecture. The above conjecture is one of the main open problems in sub-Riemannian geometry [A-openproblems, Prob. 3], [RT-MorseSard, Conj. 1]. In general, the problem is settled in the following cases:
    Thanks to the work of Agrachev-Sarychev [AS-Morse] and Agrachev-Lee [AAPL-OT], in absence of so-called Goh singular minimizing paths, all sub-Riemannian minimizing paths (resp. sub-Finsler, under suitable smoothness assumptions on the norm) are solution of a finite-dimensional Hamiltonian flow so that the corresponding set of critical values has zero measure by the finite-dimensional Sard theorem. Recently, Rifford proved that, more generally, the minimizing Sard conjecture holds for structures where all non-trivial Goh paths have Goh-rank $\leq 1$ almost everywhere, see [R-subdiff]. For Carnot groups, these results can be applied for example when the distribution is pre-medium fat, see [R-subdiff, (1.2)]. In passing, we remark that for generic (so, typically not Carnot) sub-Riemannian structures with distribution of rank $\geq 3$, there are no non-trivial singular Goh minimizing paths, see [CJT-generic, Cor. 2.5] and [AG-subanalitic, Thm. 8]. As a consequence the minimizing Sard conjecture holds true. Thanks to [R-subdiff, Cor. 1.4], the latter result extends the case of rank $\geq 2$.
    The minimizing Sard conjecture is true also for Carnot groups of step $\leq 3$, see [LDMOPV-Sardprop]. In this case, the problem is reduced (in a non-trivial way) to a finite-dimensional one. First, by noting that singular minimizing curves are non-singular in some proper subgroup, and then exploiting the fact that Carnot subgroups are parametrized by a finite-dimension manifold.
Finally, we refer to [ZZ-Rigid, BR-DukeMartinet, BFPR-StrongSardInventiones, BPR-Sardpreprint1, BPR-Sardpreprint3, BPR-Sardpreprint2, OV-codimensionCarnotSard, RT-MorseSard] for further works on various forms of the Sard conjecture on general sub-Riemannian structures.
The following result (Proposition˜52) connects this framework to the previous sections.
Proposition H (Polynomial properties of the Endpoint map).
@@
As a consequence of Proposition˜H, we can apply Theorem˜C to the study of Endpoint maps of Carnot groups. We record here a first immediate corollary.
Corollary I (Sard criterion for Endpoint maps).
@@
where $\beta_{0}=\beta_{0}(s,m)>1$ is the same constant of Theorem˜A. In particular, it holds:
@@
We can apply Corollary˜I to find large sets on which the Sard property holds for the Endpoint map of Carnot groups. This is a more functional-analytic approach to the Sard conjecture that, unlike previous ones, does not resort to reduction to finite-dimensional cases.
An interesting case in which the assumptions of Corollary˜I can be effectively checked is given by sets of controls that are “sufficiently regular”. We introduce the notation to state our result, see Section˜5.2 for additional details. Given $r>0$, and a closed interval $I\subset\mathbb{R}$, consider the set:
@@
endowed with the sup norm over the $r$-neighbourhood of $I$ in $\mathbb{C}$. More generally, let $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$ be the set of piecewise analytic controls:
@@
where $I_{j}=\inf I+\left[\frac{(j-1)|I|}{\ell},\frac{j|I|}{\ell}\right]$, for all $j=1,\dots,\ell$, which we equip with a suitable norm (see Section˜5.2). In Proposition˜53 we prove that the unit ball $K$ of $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$, with $I=[0,1]$, is compact in $H$ and its nn–width satisfies:
@@
By applying Corollary˜I we can therefore deduce the following result (see Theorem˜54).